OpenStudy (cutiecomittee123):

the arc corresponding to a central angle of 35 degrees and a radius of 10 feet is ____ feet long

OpenStudy (cutiecomittee123):

wouldnt it be s=r(theta) So s=35(10) S=350 feet

OpenStudy (muscrat123):

i have no idea sorry

OpenStudy (cutiecomittee123):

darn thanks anyways, its basic trig

OpenStudy (campbell_st):

well since the subtended angle is in degree you need to use \[Arc ~length = \frac{\theta}{360} \times 2 \times \pi \times r\] \[\theta = 35 ~~and~~ r = 10\]

OpenStudy (campbell_st):

you need to realize \[\frac{\theta}{360}\] is the fraction of the circle formed by the arc and then you need to find the circumference of the circle and multiply the answer by the fraction